OpenStudy (anonymous):

How do you do this the format is y-y1=m(X-X1) The problem: (-3,-5) (6,1) I need to put that in point slope form any help??

OpenStudy (anonymous):

First, find your slope by using the formula Y2-Y1/X2-X1

OpenStudy (anonymous):

i did that it is -6/-9

OpenStudy (anonymous):

but to negatives make a positive right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so it would be a [positive fraction.. then after that what would u do?

OpenStudy (anonymous):

Simplify 6/9 to 2/3.

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

The plug that 2/3 into the original formula: Y-Y1=m(X-X1) For Y1 and X1 values, pick either of your points: (-3,-5) or (6,1)

OpenStudy (anonymous):

It's basically like the other problem you had.

OpenStudy (anonymous):

what would be the answer though. i dont get how to plug in the answers can you shoe me one more times please/

OpenStudy (anonymous):

Let's just say you choose to use the point (6,1). You know slope is 2/3. So, Y-1=2/3(X-6)

OpenStudy (anonymous):

Can you solve from there?

OpenStudy (anonymous):

YES THANK U SO MUCH.. :) I LOVE YOU :) HEHEH

OpenStudy (anonymous):

lol no problem ^^

OpenStudy (anonymous):

HEHE :)